A share of stock with a beta of 0.76 now sells for $51. Investors expect the stock to pay a year-end dividend of $2. The T-bill rate is 3%, and the market risk premium is 7%. a. Suppose investors believe the stock will sell for $53 at year-end. Calculate the opportunity cost of capital. Is the stock a good or bad buy? What will investors do? (Do not round intermediate calculations. Round your opportunity cost of capital calculation as a whole percentage rounded to 2 decimal places.) b. At what price will the stock reach an “equilibrium” at which it is perceived as fairly priced today? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1
a) Opportunity cost of capital per CAPM = 3%+0.76*7% = 8.32%
Expected return = (2+53-51)/53 = 7.55%
As the expected return is less than the opportunity cost
investors will not invest in this stock. Those who hold
the stock will sell it.
b) Equilibrium price is that which is found out by discounting the expected cash flows with the OCC = (2+53)/1.0832 = $        50.78